Circular No. 8650

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2005 V2–V9 (SOHO) Additional near-sun comets have been found on SOHO website images

(cf. IAUC 8648), all being Kreutz sungrazers except for C/2005 V8 (which is not a member of a known group). C/2005 V2 was very faint and diffuse. C/2005 V3 was tiny, stellar in appearance, and very faint. C/2005 V4 was small and stellar in appearance in C3 images, and stellar with a hint of tail in C2 images, reaching mag 7.6 at 6.2R_{⊙} on Nov. 5.004 UT. C/2005 V5 was extremely faint and diffuse. C/2005 V6 was bright and stellar in appearance with no tail in C3 images, reaching mag 6.5 at 10.5R_{⊙} on Nov. 8.929; in C2 images, the comet showed a very faint tail reaching a length of ∼ 7′ at 5.1R_{⊙} on Nov. 9.271. C/2005 V7, a companion to C/2005 V6, was very faint and elongated. C/2005 V8 was extremely faint, very small, and stellar in appearance. C/2005 V9 appeared abruptly at ∼ 22R_{⊙} and brightened rapidly to 4th magnitude (reaching mag 3.9 at 11.1R_{⊙} on Nov. 14.221 in C3 images), developing a long, thin tail that reached a length of 40′ on Nov. 14.704 at 4.6R_{⊙} in C2 images.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 V2 Nov. 2.629 14^{h}27.^{m}4 −16^{o}14′ C2  HS 2005-Y02
C/2005 V3 3.979 14 32.5 −16 32 C2 HS 2005-Y02
C/2005 V4  4.488 14 25.7 −17 57 C3/2 TS 2005-Y03
C/2005 V5 7.788 14 49.0 −17 55 C2 RM 2005-Y03
C/2005 V6  8.346 14 37.6 −20 00 C3/2 HS 2005-Y03
C/2005 V7  9.314 14 54.4 −18 18 C2  HS 2005-Y03
C/2005 V8  9.604 15 08.6 −16 26 C2  HS 2005-Y07
C/2005 V9  13.488 15 01.7 −22 53 C3/2 HS 2005-Y07

SUPERNOVAE 2005md AND 2005mf M. Modjaz, R. Kirshner, P. Challis, and S. Blondin, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrogram (range 350–740 nm) of SN 2005md (cf. IAUC 8647), obtained by P. Berlind on Dec. 28.31 UT, shows it to be most probably a young type-II supernova. The spectrum consists of a featureless, blue continuum and is similar to an early spectrum of SN 1993J (Matheson et al. 2000, A.J. 120, 1487). A spectrum of SN 2005mf (cf. IAUC 8648), obtained by Berlind on Dec. 28.40, shows it to be a supernova of type Ic. The spectrum is similar to spectra of SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11) taken a few days before maximum.

2005 December 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green